State the number of atoms of each element on the reactant and on the product sides of the following equations:

  • a. CH4(g) + 2O2(g) àCO2(g) + 2H2O(g)
  • b. 4P(s) + 5O2(g) à P4O10(s)
  • c. 4NH3(g) + 6NO(g) à 5N2(g) + 6H2O(g)
  • d.6CO2(g) + 6H2O(l) à C3H12O6(aq) + 6O2(g)
Answer #1

Part a

CH4(g) + 2O2(g) = CO2(g) + 2H2O(g)

Reactant side

Compound name = CH4

Atoms in CH4 = C and H

Number of C atoms = 1

Number of H atoms = 4

Number of O atoms = 2*2 = 4

Product side

Compound name = CO2

Atoms in CO2 = C and O

Number of C atoms = 1

Number of O atoms = 2

Compound name = H2O

Atoms in H2O = H and O

Number of H atoms = 2*2 = 4

Number of O atoms = 2

Total number of O atoms = number of O atoms in CO2 + number of atoms in H2O

= 2 + 2 = 4

Part b

4P(s) + 5O2(g) = P4O10(s)

Reactant side

Number of P atoms = 4

Number of O atoms = 5*2 = 10

Product side

Compound name = P4O10

Atoms in P4O10 = P and O

Number of P atoms = 4

Number of O atoms = 10

Part c

4NH3(g) + 6NO(g) = 5N2(g) + 6H2O(g)

Reactant side

Compound name = NH3

Atoms in NH3 = N and H

Number of N atoms = 4

Number of H atoms = 4*3 = 12

Compound name = NO

Atoms in NO = N and O

Number of N atoms = 6

Number of O atoms = 6

Product side

Number of N atoms = 5*2 = 10

Compound name = H2O

Atoms in H2O = H and O

Number of H atoms = 6*2 = 12

Number of O atoms = 6

Part d

6CO2(g) + 6H2O(l) = C3H12O6(aq) + 6O2(g)

Reactant side

Compound name = CO2

Atoms in CO2 = C and O

Number of C atoms = 6

Number of O atoms = 6*2 = 12

Compound name = H2O

Atoms in H2O = H and O

Number of H atoms = 6*2 = 12

Number of O atoms = 6

Total number of O atoms = number of O atoms in CO2 + number of atoms in H2O

= 12 + 6 = 18

Product side

Compound name = C3H12O6

Atoms in C3H12O6 = C, H and O

Number of C atoms = 3

Number of H atoms = 12

Number of O atoms = 6

Number of O atoms in O2 = 6*2 = 12

Total number of O atoms = number of O atoms in C3H12O6 + number of atoms in O2

= 6 + 12 = 18